how many liters of Br2 will be produced from 3.71g of KBr in the following reaction?

2KBr+ F2 to 2KF+ Br2

1
Expert's answer
2021-03-10T06:58:57-0500

n(KBr) = 3.71 / 119 = 0.0312 mol

n(Br2) = 0.0312 / 2 = 0.0156 mol

m(Br2) = 160"\times"0.0156 = 2.496 g

V(Br2) = 2.496 / 3.1 = 0.805 ml

Answer: V(Br2) = 0.805 ml.